# Unsettling the Hegemony of Intention: Agonistic Image Generation

### v2

@@ Abstract.
Current image generation paradigms prioritize actualizing user intention — “see what you intend” — but often neglect the sociopolitical dimensions of this process. As these tools become cornerstones of the media landscape, it is increasingly evident that image generation is political, contributing to broader social struggles over visual meaning. This sociopolitical aspect was highlighted by the March 2024 Gemini controversy, where Gemini faced criticism for inappropriately injecting demographic diversity into user prompts. Although the developers sought to redress image generation’s sociopolitical dimension by introducing diversity “corrections,” their opaque imposition of a standard for “diversity” ultimately proved counterproductive. In this paper, we present an alternative approach: an image generation interface designed to embrace open negotiation along the sociopolitical dimensions of image creation. Grounded in the principles of agonistic pluralism (from the Greek agon, meaning struggle), our interface actively engages users with competing visual interpretations of their prompts. Through a lab study with 29 participants, we evaluate our agonistic interface on its ability to facilitate reflection — engagement with other perspectives and challenging dominant assumptions — a core principle that underpins agonistic contestation. We compare it to three existing paradigms: a standard interface emulating current image generation tools, a Gemini-style interface that produces “diverse” images, and an intention-centric interface that suggests aestheticized prompt refinements before generation. Our findings demonstrate that the agonistic interface enhances reflection across multiple measures, but also that reflection depends on users perceiving the interface as both appropriate and empowering; introducing diversity without grounding it in relevant political contexts was perceived as inauthentic. Our results suggest that diversity and user intention should not be treated as opposing values to be balanced. Instead, interfaces can productively surface and navigate tensions between competing perspectives, enabling users to engage with and evolve their intentions in meaningful ways.
@@
agonism, reflection, image generation interfaces
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Conference title.; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interaction paradigmsccs: Human-centered computing Interaction design theory, concepts and paradigms
1. Introduction: Unsettling the Hegemony of Intention
@@
What does “a Founding Father” look like? What does “Jesus” look like? What about “a migrant,” “a doctor,” “World War II”? The production and consumption of visual meanings is a sociopolitical endeavor. When people consider their mental image of “a Founding Father,” their preconceptions are in conversation with others’ interpretations about who a Founding Father is and — just as importantly — is not. Today, our collective “answers” to these questions increasingly draw from images that are generated by biased AI models and then proliferate in our digital media ecosystem (Jingnan, 2024; DiResta and Goldstein, 2024; Herrie et al., 2024). The political context of image generation was further underscored when in March 2024 Google’s Gemini model came under criticism for generating images with inaccurate racial and gender diversity — for example, depicting Nazi soldiers as people of color. Such examples show how the values embedded in image generation tools shape how we view and understand our social and political world. Despite the increasing politicization of image generation tools, the dominant paradigm behind most current AI image generation tools is the actualization of intention: they should enable anyone to see exactly what they were intending. This “intention is all you need” paradigm (Sarkar, 2024) neglects the political dimension of image generation. By inhibiting engagement with competing interpretations of user prompts, intention-centric paradigms run the risk of entrenching problematic user intentions (Cai et al., 2024).
@@
Certainly, the Gemini controversy was as a meaningful attempt to grapple with the political dimensions of image generation by implementing corrections to address a lack of diversity in image generation models. Yet, while Gemini’s diversity intervention may have challenged the “‘hegemony of intention”, it did so in a way that caused significant public backlash, inciting “an explosion of negative commentary from figures such as Elon Musk who saw it as another front in the culture wars” (Milmo and Kern, 2024). The swift pushback indicates that top-down interventions are ill-suited to the dynamic political landscape that image generation tools find themselves situated in. In short, the problem is that shallow efforts that do not involve users in discourse but simply alter their intent can come across to users as patronizing.
@@
In this work, we aim to build and evaluate an image generation interface that acknowledges the sensitive political contexts of image generation, including questions about diversity and representation, but unlike the Gemini case, encourages users to engage and reflect on their own intentions in relation to others on their own terms. Our design draws from agonistic pluralism (from the Greek agon meaning “struggle”), a political philosophy emphasizing the importance of conflict in the political realm and the productive ways it can be channeled instead of avoided. Given a user prompt, our image generation interface, Agonistic, collects and presents a range of interpretations of the prompt, prioritizing more controversial or challenging interpretations (Figure 1). Interpretations come with a justification grounded in concepts extracted from Wikipedia. Only after examining different interpretations and selecting one does a user then receive image generations.
A critical part of agonistic pluralism is reflection — the process of considering other viewpoints and challenging dominant assumptions. For this reason, we conduct a comparative lab evaluation of our interface for its ability to foster productive critical reflection. We build three other paradigmatic interfaces for comparison — Baseline: produces images generated from the user’s prompt with no further interaction; Diverse: rewrites the user’s prompt to be more diverse before image generation; and Reformulative: provides prompt reformulations to produce more aesthetic and preferable images, inspired by existing work in prompt reformulation. Twenty-nine participants used these interfaces to iteratively build and refine a collage of 10 images representing a human subject of their choice while thinking aloud.
@@
Across several measures of reflection, we find that Agonistic induces the highest reflection out of any interface (§5), followed by Reformulative, then Diverse. For instance, when participants went to add images to a collage, they described the image as an expansion of their original intention 21% of the time while using Agonistic, but only 8%, 7%, and 3% of the time for Reformulative, Diverse, and Baseline, respectively. In addition, we find that participants express diversity as a justification for adding images to their collage $48\%$ of the time when using Agonistic, but only $28\%$ under Diverse. Our findings confirm that participants are likely to reject Gemini-style interventions due to a lack of factuality and authenticity, while Agonistic more authentically expresses diversity by situating user prompts and the generated images in relevant sociopolitical discourse. We find that getting users to acknowledge and prioritize diversity requires situating diversity in authentic political controversies, a different finding from current interventions that either seek to impose diversity for its own sake or constrain diversity with factuality. Agonistic aids critical reflection by giving users a sense of control and provide contextualizing elements to help them interpret images. These results point to richer ways that researchers and developers of AI tools can address ethical questions in generative AI. Instead of making top-down interventions, our work shows that AI tools can instead create space for open-ended contestation over values like user intention, factuality, and diversity. Ultimately, our findings highlight the importance of bringing HCI work to bear on AI ethics and provide more general design elements for building reflective AI systems.
2. Background and Context
2.1. Images as socio-political entities
@@
Philosophers of meaning have argued that signs (things like words or images representing something other than themselves) can be “performed” to demonstrate membership within social categories (Butler, 1990; Hall, 1990, 1997) and invoked to effect changes in the world (Austin, 1956, 1962) — that is, signs do not just represent, but also play active roles in social life. Likewise, thinkers in history have challenged concepts of “objective” truth by calling attention to complex networks of competing narratives, power relations, and institutional practices (de Certeau, 1988; White, 1973) — what Michel Foucault calls “discourse” (Foucault, 1972) — that shape historical study. Thus, historical questions of “how things were really like” are far from straightforward because appealing to historical accuracy also requires considering ways that discourse shapes the historical record itself. Thinkers in media studies have further elaborated these insights in the specific context of visual media. Images are themselves political not only in the sense that interpretation of them is conditioned by cultural context (Barthes, 1981), but also in the sense that the act of creating or capturing an images always involves “imposing standards on [one’s] subject” (Sontag, 1977). For example, Laura Mulvey’s concept of the “male gaze” describes how the movement and position of the camera can frame women as passive objects of visual pleasure (Mulvey, 1975). These works suggest that all image production and consumption — including AI image generation — is implicated in social, political, and ethical issues: “individuals implicitly engage in ongoing struggles over visual dominance and its articulation with social formations through their involvement in image production, distribution, exposure, and consumption(Herrie et al., 2024). Our aim in this paper is to build an image generation interface which engages users in this complexity of image generation.
@@ 2.2. Agonistic Pluralism: A political framework for image generation
Given the political nature of image generation, what kind of political structures should be used to deal with this conflict? Some commentators have found the problem with the Gemini case to be that “tuning to ensure that Gemini showed a range of people failed to account for cases that should clearly not show a range(Milmo and Kern, 2024). If image generation is political, however, then there may not always be a clear division between cases that should or should not show diversity, but rather a pluralism of conflicting, sincerely held beliefs. For instance, though some reacted negatively to generated images of non-White Founding Fathers, public acclaim of the Broadway play Hamilton, which used people of color to portray Founding Fathers, implies that depicting historical figures in racially inaccurate ways is not necessarily always problematic. Consequently, we draw from democracy theory in this paper to inform our design because it offers a long tradition of scholarship aimed at answering questions of pluralism. Herrie et. al. suggest that within democracy theory, agonistic pluralism is apt for analyzing the political dimensions of image generation (Herrie et al., 2024). First developed by Chantal Mouffe, agonistic pluralism is a philosophy that sees struggles over meaning, like those in the Gemini case, as inherent and ineradicable parts of democratic society. The task for democracy then is not to eradicate such struggles, but to channel them into productive ends. Doing so requires at least two major conditions which we highlight here as relevant to our work. First, democracy should always allow for the possibility of political struggle by providing avenues for hegemonic views to be challenged. Second, Mouffe qualifies that agonistic pluralism also involves a basic level of respect for others: in contrast to purely antagonistic struggle in which opponents view each other as enemies to be destroyed, agonistic struggle requires viewing opponents legitimate adversaries “whose ideas we combat but whose right to defend those ideas we do not put into question(Mouffe, 2000).
2.3. Current values in image generation interfaces
@@
Many current works aim to help users produce “better” images without necessitating additional user interaction, invoking the value of intention actualization. Prompt reformulation methods automatically rewrite prompts to produce more aesthetic or generally preferred (Zhan et al., 2024; Xu et al., 2024; Datta et al., 2024; Chen et al., 2024) images, motivated by the goal of “convey[ing] the user’s intended ideas(Mo et al., 2024) or “preserving the original user intentions” (Hao et al., 2024) Other work builds interfaces to help users explore and iterate over prompts, invoking similar values: “‘align the user’s creative intentions with the model’s generation”  (Wang et al., 2024b); “align with [users’] intended creative output” (Brade et al., 2023). This emphasis on actualizing user intentions is described by Sarkar (2024) as the “intention is all you need” paradigm: “Contrary to the assumption that GenAI merely executes human intentions, it also shapes them […] induc[ing] ‘mechanised convergence,’ homogenising creative output, and reducing diversity in thought.” Cai et al. (2024) characterizes interfaces that overly emphasize cognitive ease of use as “easy to learn at the expense of the power and complexity necessary to do hard but valuable work or learn uncomfortable truths”. In response to these critiques, our work explores how users might question and revise their intentions, for example by reflecting upon limiting assumptions.
@@
Another notable value in image generation is diversity. Image generation models have well-documented issues with diversity, often failing to appropriately represent demographic minorities and resorting to stereotypes (Ali et al., 2024; Wang et al., 2023, 2024a; Chauhan et al., 2024; Naik and Nushi, 2023, inter alia). Approaches to address diversity issues include finetuning the model with some diversity objective (Miao et al., 2024; Esposito et al., 2023), intervening on internal model states (Friedrich et al., 2023; Zhang et al., 2023; Zameshina et al., 2023; Chuang et al., 2023), or rewriting the prompt to emphasize diversity (Microsoft, 2024). Our work takes another angle on diversity, aiming to engage users with competing visual interpretations of a subject rather than directly generating “diverse” images. We further discuss this issue in §7.1.
@@ 2.4. Agonistic design and reflection in HCI
In contrast to this intention-centric paradigm, Kate Crawford poses the question: “What might happen if we brought a model of agonism to understanding algorithms?(Crawford, 2016). She answers that agonistic design requires acknowledging that “complex, shifting negotiations are occurring between people, algorithms, and institutions, always acting in relation to each other(Crawford, 2016). These suggestions are in line with more recent literature in support of design principles that similarly seek to disrupt rather than reproduce user intention (Cai et al., 2024; Sarkar, 2024; Herrie et al., 2024). As applied to image generation, agonistic design entails that developers might avoid imposing a single standard on image generation outputs in favor of bringing controversy to users’ attention and providing “resources for more people to participate in that dissensus(Crawford, 2016).
Agonistic design also finds affinities in existing HCI work on designing for reflection, understood here as “intellectual and affective activities in which individuals engage to explore their experiences in order to lead to new understandings and appreciations(David Boud, 1985). Based on a review of HCI research about reflection, Fleck and Fitzpatrick set forth a five-level framework of reflection, in which each level is distinguished from the last by increasingly deep consideration of different views. The last stage of reflection, called critical reflection, involves challenging personal assumptions by considering a wide range of social and ethical issues (Fleck and Fitzpatrick, 2010). Halbert and Nathan further elaborate this last stage of critical reflection as the result of a two-stage process in which users first encounter a discomforting experience, and then become critically aware of their own biases by working through discomforting feelings, rather than avoiding them (Halbert and Nathan, 2015). Though comparatively little HCI research on reflection has been done for art and culture domains (Bentvelzen et al., 2022), we aim to fill that gap by designing image generation interfaces for user reflection in this paper.
3. Four Paradigmatic Interfaces for Image Generation Interactions

### v3

# Agonistic Image Generation: Unsettling the Hegemony of Intention

@@ Abstract.
Current image generation tools largely follow an intention-centric paradigm that aims to actualize user intentions but neglects the sociopolitical conversations that these intentions are embedded in. As these tools become cornerstones of the media landscape, however, it is increasingly evident that sociopolitical conflicts over visual representation are inescapable parts of image generation. For instance, in March 2024, Google’s Gemini faced criticism for inappropriately injecting demographic diversity into user prompts. Although Gemini challenged user intentions, its top-down imposition of a standard for “diversity” ultimately proved counterproductive. In this paper, we present an alternative approach: an image generation interface designed to embrace open negotiation along the sociopolitical dimensions of image creation. Grounded in agonistic pluralism (from the Greek agon, meaning struggle), our interface actively engages users with competing visual interpretations of their prompts. Through a lab study with 29 participants, we evaluate our agonistic interface on its ability to facilitate reflection, or engagement with other perspectives that challenges dominant assumptions. We compare it to three existing paradigms: a baseline interface that emulates current image generation tools, a Gemini-style interface that produces “diverse” images, and an intention-centric interface that suggests aestheticized prompt refinements. We find that the agonistic interface enhances reflection across multiple measures, but also that reflection depends on authentically grounding diversity in relevant political contexts. Our results suggest that diversity and user intention need not be treated as inherently opposing values. Instead, interfaces can productively navigate tensions between competing perspectives, enabling users to engage with and evolve their intentions in meaningful ways.
@@
agonism, reflection, diversity, factuality, image generation interfaces, human-AI interaction
@@
copyright: acmlicensedjournalyear: 2025copyright: ccconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732030isbn: 979-8-4007-1482-5/2025/06ccs: Human-centered computing Interaction paradigmsccs: Human-centered computing Interaction design theory, concepts and paradigms
One does not become enlightened by imagining figures of light, but by making the darkness conscious.” – Carl Jung
1. Introduction: The Hegemony of Intention
@@
What does “a Founding Father” look like? What does “Jesus” look like? What about “a migrant,” “a doctor,” “World War II”? Today, the “answers” to these questions increasingly draw from AI-generated images proliferating in our digital media ecosystem (Jingnan, 2024; DiResta and Goldstein, 2024; Herrie et al., 2024). The politics of image generation were underscored in March 2024 when Google’s Gemini model came under criticism for generating images with inaccurate racial and gender diversity, such as depictions of “Founding Fathers” as women of color (Milmo and Kern, 2024). While many found these historically inaccurate images of Founding Fathers to be problematic, public acclaim of works like the Broadway play Hamilton, which used actors of color to portray Founding Fathers, implies that depicting historical figures in “inaccurate” ways can nonetheless carry valuable artistic and political messages. These examples show how the production and consumption of visual meanings is a complex sociopolitical endeavor, and the values embedded in image generation tools shape how we engage with our social and political world. When people consider their mental image of “a Founding Father,” (or “Jesus”, “a migrant”, etc.) their preconceptions are embedded in dynamic ongoing conversations with others about who a Founding Father is and—just as importantly—is not. Image generation interfaces are therefore part of these complex conversations and should be designed with this in mind.
Despite the political character of image generation, the dominant paradigm behind many image generation tools remains what Sarkar (2024) calls an “intention is all you need” paradigm that prioritizes the actualization of intention—aiming for users to “see what they intend.” While this paradigm may make users happy, it also obscures the conflictual nature of visual representation—that users’ mental image of a subject may only be a part of a larger conversation with others have different mental images of that subject. When users do not engage with this discourse, their ideas may be unchallenged, detached, and polarized (Cai et al., 2024).
@@
The Gemini controversy certainly challenged this “hegemony of intention” by engaging users with political controversies over diversity and representation. However, it did so in a way that caused significant public backlash, inciting “an explosion of negative commentary from figures such as Elon Musk” that prompted Google to rescind its changes (Milmo and Kern, 2024). The swift pushback suggests that top-down interventions may be ill-suited to the dynamic political landscape of image generation. In short, the Gemini case demonstrates that interventions which do not actively engage users in discourse but rather impose on their intent can come across as patronizing infringements on user agency.
@@
In this work, we build and evaluate an image generation interface that recenters the political context of image generation by encouraging users to actively engage with a wide range of conflicting interpretations on their own terms. Our design draws from agonistic pluralism (from the Greek agon meaning “struggle”), a political philosophy emphasizing the importance of healthy democratic conflict. Given a user prompt, our interface Agonistic highlights ongoing political controversies over the user prompt by collecting and presenting a range of conflicting interpretations to the user (Figure 1). Only after engaging with different interpretations and selecting one does a user then receive image generations.
@@
A critical part of agonistic pluralism is reflection—a process that involves engaging with competing viewpoints and challenging dominant assumptions. For this reason, we conduct a comparative lab evaluation of our interface with 29 participants on its ability to foster productive critical reflection. We build three other paradigmatic interfaces for comparison: Baseline: produces images generated from the user’s prompt with no further interaction; Diverse: rewrites the user’s prompt to be more diverse before image generation; and Reformulative: provides prompt reformulations to produce more aesthetic and preferable images, inspired by existing work in prompt reformulation. Across several measures of reflection, we find that Agonistic induces the highest reflection out of any interface (§5), followed by Reformulative, then Diverse. For instance, when participants went to add images to a collage, they described the image as an expansion of their original intention 21% of the time while using Agonistic, but only 8%, 7%, and 3% of the time for Reformulative, Diverse, and Baseline, respectively. Our findings confirm that participants are likely to reject Gemini-style interventions due to a lack of factuality and authenticity, while Agonistic more authentically expresses diversity by situating the user prompt in relevant sociopolitical discourse. Thus, getting users to acknowledge and prioritize diversity requires situating diversity in authentic political controversies, departing from existing interventions that either seek to impose diversity for its own sake or constrain diversity with factuality. Instead of imposing top-down interventions onto users, AI tools can instead create space for open-ended democratic contestation over values like user intention, factuality, and diversity. Our findings highlight the importance of bringing HCI work to bear on AI ethics and provide more general design elements for building reflective AI systems.
2. Background and Related Work
2.1. Images as Sociopolitical Artifacts
@@
Philosophers of semiotics have long argued that signs (symbols like words or images representing something other than themselves) can be “performed” to demonstrate membership within social categories (Butler, 1990; Hall, 1990, 1997) and invoked to effect changes in the world (Austin, 1956, 1962). In other words, signs do not just represent, but also play active roles in social life. Likewise, philosophers of history have challenged concepts of “objective” truth by calling attention to complex networks of competing narratives, power relations, and institutional practices (de Certeau, 1988; White, 1973)—what Michel Foucault calls “discourse” (Foucault, 1972)—that shape historical record and study. Thus, historical questions of “how things were really like” are far from straightforward because appealing to historical accuracy also requires considering how discourse shapes the historical record itself. Media studies scholars have further elaborated upon these foundational insights in the context of visual media. Images are themselves political not only in the sense that interpretation of them is conditioned by cultural context (Barthes, 1981), but also in the sense that the act of creating or synthesizing images always involves “imposing standards on [one’s] subject” (Sontag, 1977). These works suggest that all image production and consumption—including AI image generation—is implicated in social, political, and ethical issues (Mulvey, 1975): “individuals implicitly engage in ongoing struggles over visual dominance and its articulation with social formations” (Herrie et al., 2024). Our aim in this paper is to build an image generation interface which engages users in this complexity of image generation.
@@ 2.2. Agonistic Pluralism: A Political Framework for Image Generation
Given the political nature of image generation, what kind of political structures should be used to deal with issues of diversity and representation? Some commentators have found the problem with the Gemini case to be that “tuning to ensure that Gemini showed a range of people failed to account for cases that should clearly not show a range(Milmo and Kern, 2024). However, scholars like Melamed (2014) argue that too often, categories of “good” and “bad” diversity only end up legitimizing existing relations of domination—“good” forms of diversity are those that comport with hegemonic social narratives, whereas “bad” forms of diversity are those that challenge them. An alternative approach is to abandon the expectation of a clear division between good and bad forms of diversity, and instead embrace a radically democratic approach that recognizes an arena of conflicting views that challenge dominant assumptions. For instance, as alluded to in §1, the differing social responses to the Gemini controversy and Hamilton can be thought of as part of a larger conflict over representations of American culture and history.
To this effect, Herrie et al. (2024) propose agonistic pluralism as a particularly apt theory of democracy for negotiating the political dimensions of image generation. First developed by Mouffe (2000), agonistic pluralism is a philosophy that sees political life as being inescapably structured by conflict between competing views. The task for democracy is thus not to eradicate such conflict, however, but to channel it into productive ends. To do so, democracy should always allow for the possibility of political conflict by providing avenues to challenge hegemonic assumptions. However, it should also work to construct political conflict in adversarial rather than antagonistic terms: in contrast to purely antagonistic struggle in which opponents view each other as enemies to be destroyed, Mouffe explains that agonistic struggle requires viewing opponents legitimate adversaries “whose ideas we combat but whose right to defend those ideas we do not put into question(Mouffe, 2000). Agonistic pluralism is therefore critical of top-down interventions like the Gemini case which impose a singular standard of visual representation and eliminate the possibility for users themselves to engage in broader political conflicts over visual prompts.
2.3. Values in Image Generation Interface Design
@@
Many current works aim to help users produce “better” images without necessitating additional user interaction, invoking the value of intention actualization. Prompt reformulation methods automatically rewrite prompts to produce more aesthetic or generally preferred (Zhan et al., 2024; Xu et al., 2024; Datta et al., 2024; Chen et al., 2024) images, motivated by “convey[ing] the user’s intended ideas(Mo et al., 2024) or “preserving the original user intentions” (Hao et al., 2024). Other work builds interfaces to help users explore and iterate over prompts, invoking similar values: “align the user’s creative intentions with the model’s generation”  (Wang et al., 2024b); “align with [users’] intended creative output” (Brade et al., 2023). This emphasis on actualizing user intentions, and its potential pitfalls, is emblematic of what Sarkar (2024) calls the “intention is all you need” paradigm. “Contrary to the assumption that GenAI merely executes human intentions, it also shapes them […] induc[ing] ‘mechanised convergence,’ homogenising creative output, and reducing diversity in thought.” Interfaces that overly emphasize ease of use may be “easy to learn at the expense of the power and complexity necessary to do hard but valuable work or learn uncomfortable truths” (Cai et al., 2024). In response to these critiques, our work explores how users might question and revise their intentions, for example by reflecting upon limiting assumptions.
@@
Another notable value in image generation is diversity, especially human diversity, which we focus on in this paper. Image generation models have well-documented issues with diversity, often failing to appropriately represent demographic minorities and resorting to stereotypes (Ali et al., 2024; Wang et al., 2023, 2024a; Chauhan et al., 2024; Naik and Nushi, 2023, inter alia). Proposed solutions to these issues include finetuning the model with some diversity objective (Miao et al., 2024; Esposito et al., 2023), intervening on internal model states (Friedrich et al., 2023; Zhang et al., 2023; Zameshina et al., 2023; Chuang et al., 2023), or rewriting the prompt to emphasize diversity (Microsoft, 2024). Our work takes a different, interactive angle on diversity, aiming to engage users with competing visual interpretations of a subject rather than directly generating “diverse” images (see §7.1).
@@ 2.4. Agonistic Design and Reflection in HCI
In contrast to the intention-centric paradigm, Crawford (2016) proposes an agonistic design paradigm which begins from the premise that “complex, shifting negotiations are occurring between people, algorithms, and institutions, always acting in relation to each other.” By foregrounding such political negotiations, agonistic design is also similar to design approaches like adversarial design (DiSalvo, 2012), antagonistic design (Cai et al., 2024), and other approaches that seek to disrupt rather than reproduce user intentions in human-computer interactions (Sarkar, 2024; Herrie et al., 2024). Agonistic design also finds affinities in HCI work on designing for reflection, where reflection is commonly understood as “intellectual and affective activities in which individuals engage to explore their experiences in order to lead to new understandings and appreciations(David Boud, 1985). Fleck and Fitzpatrick (2010) build on this definition with a five-level schema of reflection, in which higher levels of reflection involve increasingly deep engagement with other views and critical examination of biases. The highest level, critical reflection, can be understood as the result of a two-stage process in which users first encounter a discomforting experience, and then become critically aware of their own biases by working through discomforting feelings (Halbert and Nathan, 2015). Angeli et al. (2021) thus conceive of critical reflection as a key aim of agonistic pluralism and propose three tenets of agonistic design for video games to support critical reflection: confronting users with unsettling choices, encouraging engagement with multiple perspectives, and situating subject matter in a context that resonates with users. Nonetheless, art and culture remain an underexplored area of HCI research on reflection (Bentvelzen et al., 2022), and to our knowledge no prior work has applied agonistic design to image generation tools.
3. Four Image Generation Interfaces